Imagine that we chop a continuous laser beam (assumed to be monochromatic at λ0 = 632.8 nm) into tp = 0.1 ns pulses, using some sort of shutter with a chop rate 1/tc = 10 kHz.

(a) Compute the resultant laser bandwidth ∆ν, assuming tc >> tp.

(b) How does your answer change if we could chop at 1015 Hz?
Answer #1

Solution Given That Xo = 632.8 nm - 632.88109 m ste= oins = 1010 Here The Coherence length of laser beam is ДТ.: с. БІ Here C

Where w (Angular frequey)= QTU k (propageration consteut), Alto) (to) DV = V (-1/287) DV = x (Dio) - from 020 c (neglect nega

sro - (632.8x1093% 3x102 - 4.0043584 x 103 3410²2 = 1.3347X1041 Ox-0.01330m on the frequency bandwidth Av=105 HZ DI c = cх At

Please comment if you have any doubt

Please give me like because of my future contribution it's needed..Thank u :)